OpenStudy (ckallerid):

I need help .-. Factor completely 169x^2 − 64. A (13x + 8)(13x − 8) B (13x + 8)^2 C (13x − 8)^2 D 13x − 8

Nnesha (nnesha):

difference of squares \[\huge\rm a^2-b^2=(a+b)(a-b)\]

OpenStudy (rushwr):

I think it is A

OpenStudy (rushwr):

\[\sqrt{169} = 13\]

OpenStudy (rushwr):

\[169x ^{2}-64 = 13^{2}x ^{2}-8^{2}\]

OpenStudy (rushwr):

So as to what Nnesha explained it is going to be answer A

OpenStudy (ckallerid):

okie thanks! :)

OpenStudy (rushwr):

anytym @ckallerid